# Unsupervised anomaly detection in MeV ultrafast electron diffraction

### v1

@@ IV.2 Statistical analysis of the reconstruction error
Fig. 7 shows the result of the statistical analysis after this optimization, where both joint probability estimations $p(e,H)=p(H)p(e|H)$ are represented in a dash-dotted line for $H=N$ and in a dashed line for $H=A$. The optimized values for the parameters were $\nu_{N}=1.82,\mu_{N}=0,\alpha_{N}=0.17,\nu_{A}=0,\mu_{A}=2.94,\alpha_{A}=0.25,w=0.6$. Note here that for the case of faulty images (H=A), the shape parameter $\nu_{A}$ is zero, and therefore the distribution becomes a biased Rayleigh distribution. The distribution for the normal images is not biased ($\mu_{N}=0$). Finally, the prior of the normal images is $p(N)=w=0.6$, as expected since this is the fraction of normal images in the dataset.
IV.3 Estimation of the detection threshold
@@
IV.4 Experiment with reduced number of anomalies
@@
V Conclusions
We presented a methodology for anomaly detection applied to detecting of faulty images that is fully unsupervised. This is, the system is not provided with any information about what images are normal or faulty, which allows the user to train the method without the necessity of labelling the training images as normal or faulty. This way, the method can be trained and tested with a high number of images in short time.
@@ V Conclusions
The method provides the user with a an estimation of the probability of that the image is normal. This informs the user about the images that are difficult to classify for them to visually inspect a posterior. The detection threshold is computed automatically using a set of validation images that do not need to be labelled either, which makes this system fully flexed and totally autonomous, freeing the user of manually adjusting any parameters of the detector.
@@
The methodology has been trained with only 100 images. The test images were 1521, that were used first to estimate the threshold. After this, these images were tested to detect the faulty ones, with a false positive rate that ranged between 0.2 and 0.4%. The training time was of about 10 seconds per image, and the test time was of about one second per image.
Acknowledgements.
We would like to thank the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing the high performance computing, large-scale storage, and visualization resources used in this work. We wish to acknowledge the support of

### v2

@@ 4.2 Statistical analysis of the reconstruction error
Fig. 7 shows the result of the statistical analysis after this optimization, where both joint probability estimations $p(e,H)=p(H)p(e|H)$ are represented in a dash-dotted line for $H=N$ and in a dashed line for $H=A$. The optimized values for the parameters were $\nu_{N}=1.82,\mu_{N}=0,\alpha_{N}=0.17,\nu_{A}=0,\mu_{A}=2.94,\alpha_{A}=0.25,w=0.6$. Note here that for the case of faulty images (H=A), the shape parameter $\nu_{A}$ is zero, and therefore the distribution becomes a biased Rayleigh distribution. The distribution for the normal images is not biased ($\mu_{N}=0$). The prior of the normal images is $p(N)=w=0.6$, as expected since this is the fraction of normal images in the dataset.
Fig. 7 shows four images that are classified as normal, but that show errors of amplitudes 1.41, 2.03, 2.74 and 3.2. These images contain artifacts of high amplitude, which is over 1000 times. The CAE is not able to reconstruct these amplitudes, and therefore, the error highly deviates from the error mode of the normal images. The average maximum amplitude of the images is of about 10 times lower than the amplitude of these images. The method does not distinguish such subtle anomalies, and further refinements should be applied to detect these artifacts.
Additional experiments were conducted to assess the method’s discrimination performance when discrimination is performed at the tile level during testing. This approach reduces test time by approximately a factor of three, while yielding no noticeable difference in discrimination performance.
4.3 Estimation of the detection threshold
@@
4.4 Experiment with reduced number of anomalies
@@
5 Conclusions
We presented a methodology for anomaly detection applied to the detection of faulty diffraction patterns in single shot MUED experiments that is fully unsupervised. This is, the system is not provided with any information about what images are normal or faulty, which allows the user to train the method without the necessity of labeling the training images as normal or faulty. This way, the method can be trained and tested with a high number of images in short time and faulty images can be removed effectively. This can lead to increased resolution in MUED by improving the signal-to-noise ratio in averaged diffraction patterns and can also be employed as a stability diagnostic tool for the system by tracking the faulty detection rate. For pump-probe MUED experiments, where subtle changes are expected in the diffraction patterns, this methodology can dramatically increase the technique accuracy. This unsupervised filtering procedure could also be applied to aid users of other diffraction techniques (such as in-situ transmission electron diffraction) in which large datasets are collected that include faulty images due to instrumental instabilities. It should be noted, however, that this filtering procedure is not a substitute for stability and considers outliers as rare events.
@@ 5 Conclusions
The method provides the user with an estimation of the probability of that the image is normal. This informs the user about the images that are difficult to classify for them to visually inspect a posterior. The detection threshold is computed automatically using a set of validation images that do not need to be labelled either, which makes this system fully flexed and totally autonomous, freeing the user of manually adjusting any parameters of the detector.
@@
The methodology has been trained with only 100 images. The test images were 1521, which were used first to estimate the threshold. After this, these images were tested to detect the faulty ones, with a false positive rate that ranged between 0.2 and 0.4%. The training time was of about 10 seconds per image, and the test time was of about one second per image.
The core anomaly detector idea is general, but our current pre-processing is tuned to crystalline patterns dominated by localized Bragg peaks. Indeed, in order to select the image areas relevant for the training, we perform a tile selection through the estimation of connected above-threshold pixels. For applications as gas-phase diffraction the same framework can apply, but the pre-processing/tiling criterion should be adapted to that purpose.
Subtle localized defects (e.g., dark-current patches, hot pixels, partial region dropout) will increase reconstruction residuals if they are not part of the learned normal distribution. The CAE residual approach is sensitive when we aggregate errors over tiles and pixels. However, small or weak localized phenomena will be blurred if we use a global mean of the images. In these cases, it is better to use aggregations that are more robust. For example, we could use the maximum tile error, upper percentiles of tile errors, or others, while using the same probabilistic approach for the post-processing of the residuals.
Some additional work is possible to further exploit the information captured in the encoder’s latent space. Although we conducted experiments using the bottleneck output to discriminate images with a Support Vector Machine (SVM) employing a squared-exponential kernel to distinguish between normal images and anomalies, these results did not outperform the unsupervised CAE residual analysis presented in this paper. Nevertheless, it may be fruitful to use a dataset labeled by anomaly type and leverage the latent representations to train a supervised classifier for anomaly type detection.
Acknowledgments
The Work was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Science and Engineering Division, Program of Electron and Scanning Probe Microscopes, award DE-SC0021365. Funding was available through the Department of Energy’s Established Program to Stimulate Competitive Research (EPSCoR) State-National Laboratory Partnerships program in the Office of Basic Energy Sciences. This work was also partially supported by the U.S. National Science Foundation under Award PHY-1549132, the Center for Bright Beams. This research used resources of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. This research used resources of the UNM Center for Advanced Research Computing, supported in part by the National Science Foundation, for providing the high-performance computing, large-scale storage, and visualization resources used in this work. This research used resources of Brookhaven National Laboratory’s Accelerator Test Facility, an Office of Science User Facility. Manel Martínez-Ramón has been partly supported by the King Felipe VI Endowed Chair of the University of New Mexico.